OpenStudy (anonymous):

(8x3 - 125) ÷ (2x - 5) please explain step by step how to do this

OpenStudy (thomas5267):

Either do a long division: http://www.purplemath.com/modules/polydiv2.htm Or do a synthetic division: http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

did not help. can someone show me the steps here?

OpenStudy (anonymous):

(8x3=24 - 125 = (-)101) divide (2x-5= (-)10) now (-)101 divide (-)10=10.1

OpenStudy (anonymous):

welcome to openstudy :)

OpenStudy (anonymous):

thank you for the welcome. what is the answer to this problem?

OpenStudy (anonymous):

10.1

OpenStudy (anonymous):

this is Polynomial Long Division. the answer is 4xsquared+10x+25 i just don't understand the steps. can someone help?

OpenStudy (anonymous):

(8*3 - 125) ÷ (2 x - 5)